# Mechanically concealed holes

### v2

@@ Abstract
When a hole is introduced into an elastic material, it will usually act to reduce the overall mechanical stiffness. A general ambition is to investigate whether a stiff shell around the hole can act to maintain the overall mechanical properties. We consider this effect from a macroscopic continuum perspective down to atomistic scales. First, we focus on the basic continuum example situation of an isotropic, homogeneous, linearly elastic material loaded uniformly under compressive plane strain for low concentrations of holes. As we demonstrate, the thickness of the shell can be adjusted in a way to maintain the overall stiffness of the system. We derive a corresponding mathematical expression for the thickness of the shell that conceals the hole. Thus, one can work with given materials to mask the presence of the holes simply by adjusting the thickness of the surrounding shells, with no need to change the materials. Our predictions from linear elasticity continuum theory are extended to atomistic levels using molecular dynamics simulations of a model Lennard-Jones solid. These extensions attest the robustness of our predictions down to atomistic scales. Thus, they open a straightforward possibility to adjust the strategy of mechanical cloaking via atomistic manipulations. From both perspectives, the underlying concept is important in the context of light-weight construction.
@@ I Introduction
Although linear elastodynamic cloaking of cylindrical holes was discussed for infinitesimal in-plane deformations Yavari and Golgoon (2019), mechanical concealment in elastostatic situations still receives growing attention. Notably, when viewed from the perspective of optimized materials design, elastostatic cloaking has been outlined following two strategies. Either the displacement fields Fachinotti, Peralta, and Albanesi (2018) or the elastic moduli Sozio, Shojaei, and Yavari (2023) and thus the type of employed materials were used as design parameters. Interestingly, a recent analytical and numerical work Fielding (2024) showed that elastostatic mechanical cloaking of a circular inclusion in two-dimensional geometries can be attained by coating the inclusion using several concentrically arranged circular rings and tuning their shear moduli. This strategy eliminates the complexity of additionally tuning the Poisson ratio.
@@
To this end, we present an alternative strategy to illustrate the concept of mechanically concealed holes using a basic example situation, keeping the basic geometry as simple as possible. We start from continuum elasticity theory and address uniform static loading under plane-strain conditions of a material that is homogeneous, isotropic, linearly elastic, and, in principle, infinitely extended. The cylindrical holes that we introduce are of sufficiently low concentration so that we may neglect their mutual mechanical interaction. In reality, we may, for instance, think of cylindrical holes drilled into plates or blocks of material. Using analytical calculations, we demonstrate that, indeed, under such conditions, we can introduce stiff cylindrical shells around the cylindrical holes so that the presence of the holes is effectively masked and mechanically concealed. Our sole tuning parameter is the thickness of the shell surrounding the hole. The concept is then extended to atomistic scales using molecular dynamics simulations of an appropriate Lennard-Jones model solid. It maintains valid for discrete particulate structures. We mention that, in reality, colloids often provide example systems to visually bridge between atomistic and macroscopic scales and may serve for corresponding illustrations Herlach et al. (2010); Royall et al. (2024); Menath et al. (2023).
@@ II.1 Theoretical background
If the associated strain $\bm{\varepsilon}$ derives from a displacement field $\mathbf{u}$, the strain must satisfy certain compatibility conditions. They ensure that our strain can be expressed as
@@
where T marks the transpose. Specifically, in our two-dimensional plane-strain geometry, these compatibility relations reduce to
@@ II.3 Thickness of the shell for mechanical concealment
Thus, where possible, selecting most beneficial materials provides advantages. Leveraging the existence of various unconventional materials, one may even opt for materials of near-zero or negative Poisson ratios Soman et al. (2012); Jiang and Park (2014); Yu, Yan, and Ruzsinszky (2017); Liu and Zhang (2018); Lee et al. (2019); Gaal et al. (2020); Gao et al. (2024) to achieve advanced mechanical shielding. However, once a choice is made or the materials are determined by other factors, then still for many practical material combinations the strategy of simply adjusting the thickness of the shell provides an effective, realizable way for mechanical concealment.
@@ III Molecular dynamics simulations of mechanical cloaking
To proceed a step beyond continuum considerations, we seek to investigate whether our analytical solution for mechanical concealment of a hole carries over to atomistic levels. If consistency between continuum and atomistic solutions is identified, it will indicate a robust cloaking strategy for large-scale applications. Thus, a scope of microscopic realization and adjustment of atomistic mechanical properties is conceivable. This correlation would allow us to connect continuum elasticity with microscopic structural stability of matter in the context of cloaking. To this end, we analyze a microscopic model solid using molecular dynamics (MD) simulations. For given combinations of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ and $b/a$. we test whether mechanical concealment can be achieved.
@@ III.1 Simulation details and protocols
Molecular dynamics (MD) simulations are carried out in two dimensions using 46200 atoms interacting via a truncated and force-shifted Lennard-Jones (LJ) potential, see Fig. 10. It is defined as (Toxvaerd and Dyre, 2011)
$\begin{split}V_{SF}(r)=\begin{cases}V_{LJ}(r)-V_{LJ}(r_{c})\\ \quad-(r-r_{c})V_{LJ}^{\prime}(r_{c}),&r<r_{c},\\[6.0pt] 0,&r\geq r_{c},\end{cases}\end{split}$ (35)
@@
where $V_{SF}(r)$ and $V_{LJ}(r)$ are the force-shifted and the standard Lennard-Jones interatomic potentials, respectively. Here, $r$ denotes the center-to-center distance between two atoms. $r_{c}$ denotes the cutoff distance of the potential (=2.5=2.5 Å). As described in the main text, to mimic a “toy model” of a strongly cohesive two-dimensional background solid, $\epsilon=\epsilon^{\mathrm{o}}$ = 1.0 eV and $\sigma$ = 1.0 Å have been chosen as energy and length scale parameters respectively. Within the shell surrounding the hole, we set $\epsilon=\epsilon^{\mathrm{i}}$. The Lorentz-Berthelot mixing rule ($\epsilon=\sqrt{\epsilon^{\mathrm{i}}\epsilon^{\mathrm{o}}}$) is used to define the energy parameter of the potential between an atom belonging to the shell and an atom being part of the surrounded elastic solid. $\sigma$ = 1.0 Å is considered for all atoms in the solid.
@@
Holes were carved out mostly with a radius $a$ of 20 Å. However, several other hole radii or ratios of hole radius $a$ to the system size $L$ ($0.009<a/L<0.09$) are considered for a particular given ratio of the mechanical moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ in order to study the effect of $a/L$ on the compressive response of the systems. To this end, computational isotropic compression tests are performed on the equilibrated samples via unbiased area contraction using a constant engineering strain rate of $10^{-4}$ per timestep. We run these simulations for $5\times 10^{6}$ MD steps (corresponding to 0.5 ns) within NVT ensemble, out of which only the initial small-strain, linear pressure-areal strain regimes (up to 0.3 $\%$ areal strain) are used to compute the bulk modulus $K$ in two dimensions via
@@
Mechanical concealment of holes is finally achieved using the following protocol. First, we choose a given stiffness of the shells surrounding the holes using a specific ratio of shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. Then, we tune the thickness of the shell by suitably varying the ratio between the outer radii of shell and hole $b/a$. The concealment ratio is the desired $b/a$ that gives identical bulk modulus ($K$) of the solid with hole to that of the pristine solid without any hole (within a relative error of $\leq 0.07\%$). Throughout our MD simulations we consider $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ = 0.34 (see Section III.3 for details). All molecular dynamics simulations described in this article are performed using LAMMPS (Thompson et al., 2022).
@@ III.4 MD results: shell thickness for concealment
Figure 6: Atomistic distributions of instantaneous virial shear stress during isotropic compression at an areal strain of 0.3 $\%$. We consider (a) a pristine elastic solid without any hole, (b) an elastic solid with an unshielded hole ($b/a$ = 1), (c) an elastic solid with a shielded hole of parameters $b/a=1.173$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$, and (d) an elastic solid with a shielded hole and parameters $b/a=2.018$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$. The outer perimeter of the shielding shell of radius $b$ is shown via a black, dashed circle in (c) and (d).
@@
IV Concealment ratio: comparing atomistic and continuum results
Quite remarkably, as shown in Fig. 7, results from atomistic simulations match well the results from analytical continuum elasticity theory over a wide range of ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ for the considered $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34. This suggests that the picture provided by continuum elasticity theory remains valid even down to atomistic scales for the considered discrete systems of particles interacting solely via LJ pair potentials. As we gradually reduce the stiffness of the shell surrounding the hole in terms of the ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ in Fig. 7, theory and simulations follow the same overall trend. Yet, quantitative differences appear in the necessary thickness of the shell around the hole that is necessary for effective concealment, as measured by $b/a$.
@@ IV Concealment ratio: comparing atomistic and continuum results
Figure 7: Comparison between the results from continuum elasticity theory and MD simulations. For given materials of $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34, we vary the ratio of the elastic moduli between the concealing shell around the enclosed hole and the surrounding elastic solid $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. We determine the ratio between the radius of the stiff shell $b$ that is necessary to mechanically conceal the presence of the enclosed hole of radius $a$ under uniform compression. As a result, we find the same qualitative trend of $b/a$ and quantitative agreement for stiff shells, that is, at elevated $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. (In the MD simulations, we set $a=$ 20 Å, implying a ratio between $a$ and box size $L$ of $a/L=0.089$.)
One possible source for these deviations may be found in a progressively pronounced inhomogeneity on the atomic scale, in relation to the finite size of the simulation box. Indeed, the number of atoms within the shell, interacting via stronger interatomic forces compared to the background solid, grows from 0.87 $\%$ to 7.26 $\%$ of the total number of atoms in the solid with lower ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$, see Fig. 8(a). Elevated thickness of the shell is necessary to conceal the hole with softer shell materials (lower $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$). Simultaneously, for fixed $a$ (here $20$ Å), as shown in Fig. 8(b), elevated $b/a$ implies an increasing outer perimeter of the shell. This implies a larger number of particles interacting inhomogeneously, that is, with different LJ parameters concerning their interaction partners . Conversely, for stiffer and thus thinner shells (higher $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$), the results from continuum theory and atomistic simulations are almost identical.
Figure 8: (a) Number of atoms within the shell (Nshell) relative to the total number of atoms in the system (N) as a function of the ratio of shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ when the presence of the hole is effectively concealed. (b) For the same systems variation of the required outer perimeter of the shell as a function of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. Dashed black lines are guides to the eye.
We remark that further quantitative agreement between analytical results and those from MD simulations can be achieved by reducing the ratio $a/L$ between the radius of the hole and the size of the simulated system. A corresponding demonstration for $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$ is shown in Fig. 9, where we observe that this difference gradually decreases as we decrease $a/L$ from 0.089 to 0.009. We remark that, for this purpose, in our simulations we stepwise reduced the radius of the hole $a$ as indicated in the legend.
Yet, generally, if we reduce the size $a$ of the hole, the deviation in bulk elastic modulus is not that large already from the start. In Fig. 5(a) and (b) the red and black curves would be very close to each other, and so would be the blue curve for concealment. Finally, concerning actual parameter values of real materials, a hollow Al shell surrounded by a background solid of epoxy resin would have $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ within the range of 0.3–0.35, and $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}\approx 10$.
Figure 9: Ratio $b/a$ between outer radius of the shell $b$ and the radius of the hole $a$ to mechanically conceal the hole under isotropic compression. With decreasing $a$, its ratio with respect to the size of the MD system $L$ decreases as well. We chose a specific, low value of the stiffness ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$. The blue line marks the result from continuum elasticity theory.
V Conclusions

### v3

@@ Abstract
When a hole is introduced into an elastic material, it will usually act to reduce the overall mechanical stiffness. A general ambition is to investigate whether a stiff shell around the hole can act to maintain the overall mechanical properties. We consider this effect from a macroscopic continuum perspective down to atomistic scales. For this purpose, we focus on the basic continuum example situation of an isotropic, homogeneous, linearly elastic material loaded uniformly under compressive plane strain for low concentrations of holes. As we demonstrate, the thickness of the shell can be adjusted in a way to maintain the overall stiffness of the system. We derive a corresponding mathematical expression for the thickness of the shell that conceals the hole. Thus, one can work with given materials to mask the presence of the holes simply by adjusting the thickness of the surrounding shells, with no need to change the materials. Our predictions from linear elasticity continuum theory are extended to atomistic levels using molecular dynamics simulations of a model Lennard-Jones solid. These extensions attest the robustness of our predictions down to atomistic scales. Thus, they open a straightforward possibility to adjust the strategy of mechanical cloaking via atomistic manipulations. From both perspectives, the underlying concept is important in the context of light-weight construction.
@@ I Introduction
Although linear elastodynamic cloaking of cylindrical holes was discussed for infinitesimal in-plane deformations Yavari and Golgoon (2019), mechanical concealment in elastostatic situations still receives growing attention. Notably, when viewed from the perspective of optimized materials design, elastostatic cloaking has been outlined following two strategies. Either the displacement fields Fachinotti, Peralta, and Albanesi (2018) or the elastic moduli Sozio, Shojaei, and Yavari (2023) and thus the type of employed materials were used as design parameters. We mention that solids containing hollow cylindrical shells, at least concerning mathematical approaches, can be viewed as one representative example of composite materials. Over time, they were studied extensively, with the seminal works of Eshelby marking a notable point of reference Eshelby (1957). Others followed Christensen (1993, 2005), where studies on composites were concerned mostly with their effective mechanical properties Christensen and Lo (1979). Analytical works usually rely on linear elasticity Hsiao-Sheng and Acrivos (1978), and generalized self consistent schemes have been employed Benveniste (2008). Specific progress on the particle scale considering mutual interactions in analytical theory was achieved when perfectly rigid inclusions were considered Phan-Thien and Kim (1994); Puljiz and Menzel (2019). Composite structures with holes have also been studied to alleviate stress around holes, albeit using functionally graded materials with varying elastic constants, while keeping Poisson ratios fixed Sburlati (2013); Sburlati, Atashipour, and Atashipour (2014). Interestingly, a recent analytical and numerical work Fielding (2024) showed that elastostatic mechanical cloaking of a circular inclusion in two-dimensional geometries can be attained by coating the inclusion using several concentrically arranged circular rings and tuning their shear moduli. This strategy eliminates the complexity of additionally tuning the Poisson ratio.
@@
Our goal is to present an alternative strategy of mechanical concealment that extends from a macroscopic, continuum level down to microscopic, atomistic scales. To illustrate this concept of mechanically concealed holes, we focus on a basic example situation, keeping the basic geometry as simple as possible. We start from continuum elasticity theory and address uniform static loading under plane-strain conditions of a material that is homogeneous, isotropic, linearly elastic, and, in principle, infinitely extended. The cylindrical holes that we introduce are of sufficiently low concentration so that we may neglect their mutual mechanical interaction. In reality, we may, for instance, think of cylindrical holes drilled into plates or blocks of material. Using analytical calculations, we demonstrate that, indeed, under such conditions, we can introduce stiff cylindrical shells around the cylindrical holes so that the presence of the holes is effectively masked and mechanically concealed. Our sole tuning parameter is the thickness of the shell surrounding the hole. The concept is then extended to atomistic scales using molecular dynamics simulations of an appropriate Lennard-Jones model solid. It maintains valid for discrete particulate structures. We mention that, in reality, colloids often provide example systems to visually bridge between atomistic and macroscopic scales and may serve for corresponding illustrations Herlach et al. (2010); Royall et al. (2024); Menath et al. (2023).
@@ II.1 Theoretical background
If the associated strain $\bm{\varepsilon}$ derives from a displacement field $\mathbf{u}$, the strain must satisfy certain compatibility conditions. They ensure that, within the framework of linear elasticity, our strain can be expressed as
@@
where T marks the transpose, independently of whether we start from the Lagrange or Euler framework Chaikin and Lubensky (1995). Specifically, in our two-dimensional plane-strain geometry, these compatibility relations reduce to
@@ II.3 Thickness of the shell for mechanical concealment
Thus, where possible, selecting most beneficial materials provides advantages. Leveraging the existence of various unconventional materials, one may even opt for materials of near-zero or negative Poisson ratios Soman et al. (2012); Jiang and Park (2014); Yu, Yan, and Ruzsinszky (2017); Liu and Zhang (2018); Lee et al. (2019); Gaal et al. (2020); Gao et al. (2024) to achieve advanced mechanical shielding. For instance, graphene monolayers can be rendered auxetic by including holes and patterned defects Grima et al. (2015, 2018). However, once a choice is made or the materials are determined by other factors, then still for many practical material combinations the strategy of simply adjusting the thickness of the shell provides an effective, realizable way for mechanical concealment. We recall that we assumed uniform imposed deformations in the presence of one hole. Confining ourselves to linear elasticity theory, the solution for several holes can simply be superimposed. However, we neglect in this framework their mutual interactions by surrounding deformations induced by the holes. Therefore, our approach is currently confined to low densities of noninteracting holes to remain quantitative.
@@ III Molecular dynamics simulations of mechanical cloaking
To proceed a step beyond continuum considerations, we seek to investigate whether our analytical solution for mechanical concealment of a hole carries over to atomistic levels. If consistency between continuum and atomistic solutions is identified, it will indicate a robust cloaking strategy for large-scale applications. Thus, a scope of microscopic realization and adjustment of atomistic mechanical properties is conceivable. This correlation would allow us to connect continuum elasticity with microscopic structural stability of matter in the context of cloaking. To this end, we analyze a microscopic model solid using molecular dynamics (MD) simulations. For given combinations of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ and $b/a$, we test whether mechanical concealment can be achieved.
@@ III.1 Simulation details and protocols
Molecular dynamics (MD) simulations are carried out in two dimensions using 46200 atoms interacting via a truncated and force-shifted Lennard-Jones (LJ) potential, see Fig. 8 in Appendix. It is defined as (Toxvaerd and Dyre, 2011)
$split V_{SF}(r)=\cases{V}_{LJ}(r)-V_{LJ}(r_{c})\\ \quad-(r-r_{c})V_{LJ}^{\prime}(r_{c}),&r<r_{c},\\ 0,&r\geq r_{c},$ (35)
@@
where $V_{SF}(r)$ and $V_{LJ}(r)$ are the force-shifted and the standard Lennard-Jones interatomic potentials, respectively. Here, $r$ denotes the center-to-center distance between two atoms. $r_{c}$ denotes the cutoff distance of the potential (=2.5=2.5 Å). To mimic a “toy model” of a strongly cohesive two-dimensional background solid, $\epsilon=\epsilon^{\mathrm{o}}$ = 1.0 eV and $\sigma$ = 1.0 Å have been chosen as energy and length scale parameters respectively. Within the shell surrounding the hole, we set $\epsilon=\epsilon^{\mathrm{i}}$. The Lorentz-Berthelot mixing rule ($\epsilon=\sqrt{\epsilon^{\mathrm{i}}\epsilon^{\mathrm{o}}}$) is used to define the energy parameter of the potential between an atom belonging to the shell and an atom being part of the surrounded elastic solid. $\sigma$ = 1.0 Å is considered for all atoms in the solid.
@@
Holes were carved out with a radius $a$ of 20 Å. To this end, computational isotropic compression tests are performed on the equilibrated samples via unbiased area contraction using a constant engineering strain rate of $10^{-4}$ per timestep. We run these simulations for $5\times 10^{6}$ MD steps (corresponding to 0.5 ns) within NVT ensemble, out of which only the initial small-strain, linear pressure-areal strain regimes (up to 0.3 $\%$ areal strain) are used to compute the bulk modulus $K$ in two dimensions via
@@
Mechanical concealment of holes is finally achieved using the following protocol. First, we choose a given stiffness of the shells surrounding the holes using a specific ratio of shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. Then, we tune the thickness of the shell by suitably varying the ratio between the outer radii of shell and hole $b/a$. The concealment ratio is the desired $b/a$ that gives identical bulk modulus ($K$) of the solid with hole to that of the pristine solid without any hole (within a relative error of $\leq 0.07\%$). Throughout our MD simulations we consider $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ = 0.34 (see Section III.3 for details). All molecular dynamics simulations described in this article are performed using LAMMPS (Thompson et al., 2022; Lam, ).
@@ III.3 Poisson ratio evaluation
We note here that the Poisson ratio in two-dimensional hexagonal systems experiences an upper limit of 1 KW (2003); Gao et al. (2021), compared to the three-dimensional case, where it is bound by $1/2$. Our obtained Poisson ratio for the two-dimensional LJ solid was evaluated for a genuinely two-dimensional system. It is approximately 0.34, so we are well within the given limits. Comparison between the results suggested by continuum elasticity theory for the considered range of thermodynamically stable Poisson ratios under plane-strain conditions and the selected atomistic two-dimensional LJ model solid is therefore possible. The Poisson ratios listed in the expressions derived from continuum elasticity theory, Eq. (34), are still the three-dimensional ones. Therefore, to compare our MD results with continuum theory, we employ the conversion formula for plane strain Eischen and Torquato (1993); Meille and Garboczi (2001), $\nu^{\{i,o\}}=\nu^{\{i,o\},MD}/(1+\nu^{\{i,o\},MD})$. In this expression, $\nu^{\{i,o\},MD}$ is our value of approximately $0.34$ obtained from MD simulations, which converts to approximately $\nu^{\{i,o\}}\approx 0.254$ for comparison with the result in Eq. (34). The shear moduli remain identical under this conversion Eischen and Torquato (1993).
Figure 6: Atomistic distributions of instantaneous virial shear stress during isotropic compression at an areal strain of 0.3 $\%$. We consider (a) a pristine elastic solid without any hole, (b) an elastic solid with an unshielded hole ($b/a$ = 1), (c) an elastic solid with a shielded hole of parameters $b/a=1.173$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$, and (d) an elastic solid with a shielded hole and parameters $b/a=2.018$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$. The outer perimeter of the shielding shell of radius $b$ is shown via a black, dashed circle in (c) and (d). OVITO Stukowski (2009) is used for visualization.
@@ III.4 MD results: shell thickness for concealment
III.5 Comparing atomistic with continuum results
@@
Figure 7: Comparison between the results from continuum elasticity theory and MD simulations. For given materials of $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34 in two-dimensional MD simulations, we vary the ratio of the elastic moduli between the concealing shell around the enclosed hole and the surrounding elastic solid $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. We determine the ratio between the radius of the stiff shell $b$ that is necessary to mechanically conceal the presence of the enclosed hole of radius $a$ under uniform compression. For comparison, the corresponding line obtained via Eq. (34) from continuum elasticity theory is included. There, under plane-strain conditions, the effective corresponding three-dimensional Poisson ratios are given by $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ $\approx 0.254$. The results from the two different approaches, continuum theory and discrete microscopic simulations, agree well with each other. (In the MD simulations, we set $a=$ 20 Å, implying a ratio between $a$ and box size $L$ of $a/L=0.089$.)
Quite remarkably, as shown in Fig. 7, results from atomistic simulations match well the results from analytical continuum elasticity theory over a wide range of ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ for the considered $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34. This value obtained from the two-dimensional MD simulations converts into $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ $\approx 0.254$ for comparison with the theory, as mentioned above. The agreement in Fig. 7 suggests that the picture provided by continuum elasticity theory remains valid even down to atomistic scales for the considered discrete systems of particles interacting solely via LJ pair potentials. It confirms our goal of demonstrating a way of mechanical concealment under the given loading conditions that is accessible both on a macroscopic continuum level and down to microscopic, atomistic scales.
IV Conclusions